Mr. Wilson invested money in two accounts. His total investment was $30,000. If one account pays 4% in interest and the other pays 6% in interest, how much does he have in each account if he earned a total of $1,380 in interest in 1 year?
Answer by mananth(16949)

Part I 4.00% per annum ------------- Amount invested =x
Part II 6.00% per annum ------------ Amount invested = y
30000
Interest----- 1380.00

Part I 4.00% per annum ---x
Part II 6.00% per annum ---y
Total investment
x + 1 y= 30000 -------------1
Interest on both investments
4.00% x + 6.00% y= 1380
Multiply by 100
4 x + 6 y= 138000.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -120000.00
Add this to (2)
0 x 2 y= 18000
divide by 2
y = 9000
Part I 4.00% $ 21000
Part II 6.00% $ 9000

CHECK
21000 --------- 4.00% ------- 840.00
9000 ------------- 6.00% ------- 540.00
Total -------------------- 1380.00

m.ananth@hotmail.ca